# OSCAR: One-Step Diffusion Codec for Image Compression Across Multiple Bit-rates

### v3

@@ 4 Experiments
Refer to caption (c) Quantitative comparison on the CLIC2020 dataset
@@ 4.1 Experimental Settings
Datasets. We curate our training dataset from DF2K, which integrates 800 images from DIV2K [2], 2,650 from Flickr2K [45], and images from LSDIR [31], resulting in a total of 88,441 high-quality samples. For evaluation, we benchmark OSCAR against representative baselines on three standard datasets: Kodak [17], DIV2K-val [2], and CLIC2020 [46]. The Kodak dataset includes 24 natural images at a resolution of 768$\times$512, while DIV2K-val and CLIC2020 contain 100 and 428 high-definition images, respectively. All evaluation images from DIV2K-val and CLIC2020 are center-cropped to 1024$\times$1024 resolution.
Evaluation metrics. To ensure comprehensive evaluation, we adopt a set of representative full- and no-reference image quality metrics. To assess perceptual similarity with ground-truth images, we employ LPIPS [57] and DISTS [15], both of which leverage deep neural network features to capture structural and perceptual alignment. Furthermore, to evaluate visual quality in a reference-free setting, we adopt MUSIQ [26] and CLIPIQA [52], which are designed to estimate human-perceived quality directly from the input image.
Implementation details. Our approach builds upon Stable Diffusion 2.1 [39], which comprises approximately 965.9M parameters. Each bit-rate-specific hyper-encoder contains between 3.8M and 7.1M parameters. During the first training phase (Section 3.3), we optimize all hyper-encoders in parallel using the Adam optimizer [27], with a fixed learning rate of 2$\times$10-5 and a batch size of 64. Training is conducted for 10,000 iterations on a single NVIDIA RTX A6000 GPU.
In the second stage (Section 3.4), we train our model with the AdamW optimizer [33], setting the learning rate to 5$\times$10-5, weight decay to 1$\times$10-5, and batch size to 32 for both OSCAR and the discriminator. The discriminator architecture incorporates a pre-trained UNet encoder from Stable Diffusion and a lightweight MLP head. We apply LoRA [24] with a rank of 16 for efficient fine-tuning. The quantized latents produced by hyper-encoders are used as inputs corresponding to different compression levels. The discriminator follows the same training protocol as OSCAR. The perceptual loss weight $\lambda_{1}$ is set to 1, and the adversarial loss weight $\lambda_{2}$ is 5$\times$10-3. This stage is trained for 100,000 iterations using four NVIDIA RTX A6000 GPUs.
@@ 4.2 Main Results
Compared methods. We benchmark OSCAR against four representative categories of image codecs: 1) Traditional methods: BPG [6] and VVC [10], which serve as hand-crafted baselines. 2) Learning-based single-rate codecs: HiFiC [35] uses conditional GANs to balance rate, distortion, and perceptual quality, while MS-ILLM [37] enhances local realism via spatially-aware adversarial training. 3) Learning-based multi-rate codecs: DPICT [29] and CTC [25] both adopt tri-plane coding for flexible rate control, with CTC further leveraging context modeling. 4) Diffusion-based codecs: CDC [54] and DiffEIC [32] are optimized for the rate-distortion objective, while PerCo [11] adopts vector quantization. Both DiffEIC and PerCo build upon the same pretrained Stable Diffusion 2.1 model as ours. Each requires a separate model per bit-rate. For HiFiC and CDC, we additionally retrain them on our dataset at lower bit-rates to ensure fair and consistent comparison.
@@
Qualitative results. As shown in Fig. 6, traditional, learning-based, and diffusion-based codecs generally preserve the coarse structure of the image but suffer from noticeable quality degradation, such as blurring, texture loss, color shifts, and visible artifacts. For instance, wall textures often appear overly smooth, while fine structures such as window frames, door edges, and decorative patterns are frequently distorted or lost. In contrast, OSCAR preserves such details with high fidelity, maintaining both material realism and structural consistency with the original image. Overall, it achieves a highly favorable rate-distortion trade-off, ensuring that compressed outputs retain both excellent perceptual quality and faithful structural integrity.
@@ 4.3 Ablation Studies
Figure 8: Effect of representation alignment on residuals and cosine similarity. The purple curve in the left plot is the standard Gaussian.
Representation alignment. Representation alignment is key to modeling quantized latents as intermediate diffusion states, as it encourages their residuals to follow a standard Gaussian distribution. It also helps preserve structural information, enabling accurate one-step reconstruction. To validate this, we compare the residual term in Eq. (11) with and without alignment. As shown in Fig. 8, the aligned model produces residuals that closely match the standard Gaussian, whereas the unaligned one shows large deviations. Cosine similarity between quantized and original latents is also notably higher with alignment training. Figure 7 further demonstrates the effectiveness of alignment through both visual and quantitative comparisons, showing that removing it leads to substantial performance degradation.

### v6

@@ 4 Experiments
(c) Quantitative comparison on the CLIC2020 dataset
@@ 4.1 Experimental Settings
[…]
@@ 4.2 Main Results
Compared methods. We benchmark OSCAR against four representative categories of image codecs: 1) Traditional methods: BPG bpg and VVC bross2021overview , which serve as hand-crafted baselines. 2) Learning-based single-rate codecs: HiFiC mentzer2020high uses conditional GANs to balance rate, distortion, and perceptual quality, while MS-ILLM muckley2023improving enhances local realism via spatially-aware adversarial training. 3) Learning-based multi-rate codecs: DPICT lee2022dpict and CTC jeon2023context both adopt tri-plane coding for flexible rate control. 4) Diffusion-based codecs: CDC yang2023lossy and DiffEIC li2024towards are optimized for the rate-distortion objective, while PerCo careil2024towards adopts vector quantization. Both DiffEIC and PerCo build upon the same pretrained Stable Diffusion 2.1 model as ours. We retrain PerCo using the open source implementation PerCo (SD) korber2024perco . Each requires a separate model per bit-rate. For HiFiC and CDC, we additionally retrain them on our dataset at lower bit-rates.
@@
Discussion of the trend of RD-curves. The RD-curves of OSCAR are not strictly convex, as the bitrate is jointly determined by both the downsampling ratio of the hyper-encoder and the size of the codebook. When the bitrate changes, the effects introduced by these two factors may not vary consistently. Nevertheless, our method maintains smooth transitions between adjacent bitrates without noticeable performance drops, while still achieving strong results at low bitrates.
Qualitative results. As shown in Fig. 6, traditional, learning-based, and diffusion-based codecs generally preserve the coarse structure of the image but suffer from noticeable quality degradation, such as blurring, texture loss, color shifts, and visible artifacts. For instance, wall textures often appear overly smooth, while fine structures such as window frames, door edges, and decorative patterns are frequently distorted or lost. In contrast, OSCAR preserves such details with high fidelity, maintaining both material realism and structural consistency with the original image. Overall, it achieves a highly favorable rate-distortion trade-off, ensuring that compressed outputs retain both excellent perceptual quality and faithful structural integrity. We provide more results in the appendix.
@@ 4.3 Ablation Studies
(a) Effect of representation alignment. The purple curve in the left plot is the standard Gaussian.
(b) Results of generalization to unseen bitrates. The star denotes the newly adapted bitrate.
Figure 8: Ablation study on RePA and generalization to unseen bitrates.
Representation alignment. To validate the effectiveness of representation alignment, we compare the residual term in Eq. (11) with and without alignment. As shown in Fig. 8(a), the aligned model produces residuals that closely match the standard Gaussian, whereas the unaligned one shows large deviations. Cosine similarity between quantized and original latents is also notably higher with alignment training. Figure 7(b) further demonstrates the effectiveness of alignment through both visual and quantitative comparisons, showing that removing it leads to substantial performance degradation.
@@
Generalization to unseen bitrates. For each predefined bitrate, OSCAR employs an individual hyper-encoder for compression, resulting in a set of fixed-rate models. Nevertheless, we demonstrate that OSCAR exhibits strong generalization capability to unseen bitrates. Specifically, for a new bitrate, we continue training for 15k iterations under the same settings as before, except that in each iteration, the bitrate is sampled from the new one with a probability of 0.5 and from the existing set with a probability of 0.5. As shown in Fig. 8(b), the new bitrate can be effectively learned with only minor performance degradation on the previously trained bitrates.
Figure 9: PSNR and FID results on Kodak, DIV2K-val, and CLIC2020 datasets. The yellow dashed line indicates reconstruction by VAE encoding and decoding without compression.
PSNR and FID results. We present the PSNR and FID performance of state-of-the-art codecs in Fig. 9. In addition to the compared baselines, we further include IPIC xu2024idempotence and DDCM ohayon2025compressed , whose results on PSNR and FID are highly competitive. For FID evaluation, we report results only on the CLIC2020 dataset, as the other two datasets contain too few images for statistically stable estimation.
Diffusion-based codecs generally underperform traditional methods in terms of PSNR. This is mainly due to the inherent limitation of VAEs, which struggle to achieve pixel-wise accurate reconstruction. As indicated by the yellow dashed line in the figure, the reconstruction from a VAE (without compression) can even yield lower PSNR than traditional codecs at higher bitrates. Despite these challenges, OSCAR demonstrates competitive performance among multi-step diffusion-based codecs.
@@ 5 Conclusion
Acknowledgments
This work was supported by Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and the Fundamental Research Funds for the Central Universities.